OpenStudy (anonymous):

What is the solution set of |x + 5| > 1?

OpenStudy (anonymous):

x < –6 or x > 6 x < –6 or x > –4 x < –4 or x > –4 x < –4 or x > 6

OpenStudy (anonymous):

Also, What is the slope of the line which passes through (2, 6) and (0, 2)? 2 –3 4 –2

OpenStudy (anonymous):

Lastly, What is the equation of a vertical line passing through (1, 7)? y = 7 x = 1 y = 1 x = 7

OpenStudy (anonymous):

My guesses are A, C, A

jigglypuff314 (jigglypuff314):

no, no, and no -.-" |x+5| > 1 x + 5 > 1 x + 5 < -1 solve for x \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } = \frac{ (2) - (-7) }{ (0)-(2) }\] vertical line is a x = something (x , y) from (1, 7) ^ x = 1

OpenStudy (anonymous):

B, either a or d, B

jigglypuff314 (jigglypuff314):

yes, A, and yes :)